OpenStudy (anonymous):

Lateral Area Formula of a cone? Grace is making party hats for her upcoming birthday party. A single hat has a diameter of 15 centimeters and a slant height of 28 centimeters. What is the lateral area of each party hat? 836.45 cm2 659.73 cm2 2,026.33 cm2 1,319.47 cm2 *lateral area confuses me can you please explain?*

OpenStudy (anonymous):

surface area of a cone =pi*r*l r=radius =7.5cm l=slant height...=28 cm 3.14*7*28=659...